OpenStudy (anonymous):

Write 7/20 as a percent and a decimal.

OpenStudy (anonymous):

turn it into a decimal first... 7 divided by 20..then take that # move it 2 decimal places to the right and you got your answer

OpenStudy (anonymous):

035.

OpenStudy (anonymous):

@ ~Jane_Doe~

OpenStudy (anonymous):

@~Jane_Doe~

OpenStudy (anonymous):

7/20=.35=your decimal your % would be 35%

OpenStudy (anonymous):

so my decimal is .35 and 35%

OpenStudy (anonymous):

@~Jane_Doe~

OpenStudy (anonymous):

decimal=.35 percent=35%

OpenStudy (gorv):

for %...7/20 * 100=7*5=35% decimal =7/20=0.35